# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v1

@@ A.10 Augmentation strategy visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sampled 5,000 instructions obtained through these methods and randomly sampled 10,000 in-domain instructions as the background, which represents the distribution of specific domains in the public dataset. We then visualized the results using t-SNE(van der Maaten & Hinton (2008)), as shown in Figure LABEL:fig:augmentation_strategy_visualization. The plot shows that FedDCA encompasses the most of in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.

### v3

@@ A.10 Augmentation strategy visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset. We then visualized the results using t-SNE (van der Maaten & Hinton, 2008), as shown in Figure LABEL:fig:augmentation_strategy_visualization. The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.